# Learning solution operators of PDEs defined on varying domains via MIONet

### 23 Feb 2024 v1

@@ Abstract
In this work, we propose a method to learn the solution operators of PDEs defined on varying domains via MIONet, and theoretically justify this method. We first extend the approximation theory of MIONet to further deal with metric spaces, establishing that MIONet can approximate mappings with multiple inputs in metric spaces. Subsequently, we construct a set consisting of some appropriate regions and provide a metric on this set thus make it a metric space, which satisfies the approximation condition of MIONet. Building upon the theoretical foundation, we are able to learn the solution mapping of a PDE with all the parameters varying, including the parameters of the differential operator, the right-hand side term, the boundary condition, as well as the domain. Without loss of generality, we for example perform the experiments for 2-d Poisson equations, where the domains and the right-hand side terms are varying. The results provide insights into the performance of this method across convex polygons, polar regions with smooth boundary, and predictions for different levels of discretization on one task. Reasonably, we point out that this is a meshless method, hence can be flexibly used as a general solver for a type of PDE.
@@ 2 Problem setup
Since the difficulty lies in dealing with the varying domains, without loss of generality, here we consider the simplified case
@@
The case of (4-5) will be of no difficulty as long as this simplified case is solved.
@@ 3.3 Method
Note that $\sigma(f_{\Omega_{i}}^{i})$ and $\pi_{2}\circ\sigma(u_{\Omega_{i}}^{i})$ are preprocessed based on the dataset before training.
@@ 4 Numerical experiments
Quadrilateral Pentagon Hexagon Polar regions
@@ 5 Conclusions
In this paper, we introduce a framework rooted in MIONet to address PDEs with varying domains. We first broaden the approximation theorem of MIONet to encompass metric spaces. Consequently, this extension enables its application to the metric space comprising an uncountable set of polar regions. We then present an algorithm tailored for PDEs with varying domains, hence we are able to learn the solution mapping of a PDE with all the parameters varying, including the parameters of the differential operator, the right-hand side term, the boundary condition, as well as the domain. We for example perform the experiments for 2-d Poisson equations, where the domains and the right-hand side terms are varying. The results provide insights into the method’s performance across convex polygons, polar regions with smooth boundary, and predictions for different levels of discretization on one task.

### 17 Mar 2024 v2

@@ Abstract
In this work, we propose a method to learn the solution operators of PDEs defined on varying domains via MIONet, and theoretically justify this method. We first extend the approximation theory of MIONet to further deal with metric spaces, establishing that MIONet can approximate mappings with multiple inputs in metric spaces. Subsequently, we construct a set consisting of some appropriate regions and provide a metric on this set thus make it a metric space, which satisfies the approximation condition of MIONet. Building upon the theoretical foundation, we are able to learn the solution mapping of a PDE with all the parameters varying, including the parameters of the differential operator, the right-hand side term, the boundary condition, as well as the domain. Without loss of generality, we for example perform the experiments for 2-d Poisson equations, where the domains and the right-hand side terms are varying. The results provide insights into the performance of this method across convex polygons, polar regions with smooth boundary, and predictions for different levels of discretization on one task. We also show the additional result of the fully-parameterized case in the appendix for interested readers. Reasonably, we point out that this is a meshless method, hence can be flexibly used as a general solver for a type of PDE.
@@ 2 Problem setup
Since the difficulty lies in dealing with the varying domains, in order to facilitate the readers to understand, here we consider the simplified case
@@
The case of (4-5) will be of no difficulty as long as this simplified case is solved. Readers can refer to Appendix A for details of the fully-parameterized case.
@@ 3.3 Method
Note that $\sigma(f_{\Omega_{i}}^{i})$ and $\pi_{2}\circ\sigma(u_{\Omega_{i}}^{i})$ are preprocessed based on the dataset before training. An illustration of this method is shown in Figure 3. The method of the fully-parameterized case can be found in Appendix A.
Figure 3: An illustration of the method.
@@ 4 Numerical experiments
Quadrilateral Pentagon Hexagon Smooth boundary
@@
The numerical result of the fully-parameterized case is shown in Appendix A.
@@ 5 Conclusions
In this paper, we introduce a framework rooted in MIONet to address PDEs with varying domains. We first broaden the approximation theorem of MIONet to encompass metric spaces. Consequently, this extension enables its application to the metric space comprising an uncountable set of polar regions. We then present an algorithm tailored for PDEs with varying domains, hence we are able to learn the solution mapping of a PDE with all the parameters varying, including the parameters of the differential operator, the right-hand side term, the boundary condition, as well as the domain. We for example perform the experiments for 2-d Poisson equations, where the domains and the right-hand side terms are varying. The results provide insights into the method’s performance across convex polygons, polar regions with smooth boundary, and predictions for different levels of discretization on one task. We also show the additional result of the fully-parameterized case in the appendix for interested readers.
@@ Acknowledgments
Appendix
Appendix A Fully-parameterized solution operator
As we stated before, there is no difficulty in dealing with the fully-parameterized case as long as the simplified case of varying $f_{\Omega}$ is solved. Consequently we consider the fully-parameterized Poisson equation:
where all the parameters including $k$, $f$, $g$ and $\Omega$ are changing, hence the solution mapping we aim to learn is
$\mathcal{G}:(k_{\Omega},f_{\Omega},g_{\partial\Omega})\mapsto u_{\Omega}.$ (46)
Similarly, $\mathcal{G}$ can be rewritten as
$\mathcal{G}=\sigma_{2}\circ(\pi_{1},\hat{\mathcal{G}})\circ\sigma_{1},$ (47)
where
$\begin{cases}\sigma_{1}:(k_{\Omega},f_{\Omega},g_{\partial\Omega})\mapsto\left(\Omega,k_{\Omega}\circ\alpha_{\Omega}^{-1},\left(f_{\Omega}\circ\alpha_{\Omega}^{-1},g_{\partial\Omega}\circ(\alpha_{\Omega}^{-1})|_{\partial B(0,1)}\right)\right),\\ \sigma_{2}:(\Omega,u_{\Omega}\circ\alpha_{\Omega}^{-1})\mapsto u_{\Omega},\\ \pi_{1}:\left(\Omega,k_{\Omega}\circ\alpha_{\Omega}^{-1},\left(f_{\Omega}\circ\alpha_{\Omega}^{-1},g_{\partial\Omega}\circ(\alpha_{\Omega}^{-1})|_{\partial B(0,1)}\right)\right)\mapsto\Omega,\\ \pi_{2}:(\Omega,u_{\Omega}\circ\alpha_{\Omega}^{-1})\mapsto u_{\Omega}\circ\alpha_{\Omega}^{-1},\end{cases}$ (48)
and
$\begin{split}&\Omega~{}~{}~{}~{}~{}~{}~{}~{}k_{\Omega}\circ\alpha_{\Omega}^{-1}~{}~{}~{}~{}~{}~{}(f_{\Omega}\circ\alpha_{\Omega}^{-1},g_{\partial\Omega}\circ(\alpha_{\Omega}^{-1})|_{\partial B(0,1)})\mapsto u_{\Omega}\circ\alpha_{\Omega}^{-1}\\ \hat{\mathcal{G}}:=\pi_{2}\circ\sigma_{2}^{-1}\circ\mathcal{G}\circ\sigma_{1}^{-1}:\quad&K_{1}\quad\times\quad K_{2}\quad\quad\quad\quad\times\quad\quad\quad\quad K_{3}\quad\quad\longrightarrow\quad\quad C(B(0,1)).\\ &\cap~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\cap~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\cap\\ &U~{}~{}~{}~{}~{}~{}C(B(0,1))~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}C(B(0,1))\times C(\partial B(0,1))\end{split}$ (49)
Note that here we put the $f$ and $g$ together as a Cartesian product since the solution operator of Eq. (45) is linear with respect to $(f,g)$. So that we use a MIONet to learn $\hat{\mathcal{G}}$, which has three branch nets. The first branch net encodes $\Omega$, the second encodes $k_{\Omega}\circ\alpha_{\Omega}^{-1}$, and the third encodes $(f_{\Omega}\circ\alpha_{\Omega}^{-1},g_{\partial\Omega}\circ(\alpha_{\Omega}^{-1})|_{\partial B(0,1)})$. Moreover, the third branch net is set to linear. With a trained MIONet $\mathcal{M}$, we make prediction given $(k_{\Omega},f_{\Omega},g_{\partial\Omega})$ by
$u_{\Omega}^{\rm pred}=\sigma_{2}\circ(\pi_{1},\mathcal{M})\circ\sigma_{1}(k_{\Omega},f_{\Omega},g_{\partial\Omega}).$ (50)
For experiment, we generate 4500 polar regions with smooth boundary, subsequently create 4500 corresponding triangular meshes based on these regions. Utilizing these meshes, we generate 4500 random functions $f_{\Omega}$, $k_{\Omega}$, and $g_{\partial\Omega}$ for the regions, and then solve equations (45) employing the finite element method. Similar to the preceding experiments, we generate 5000 random but fixed points within $B(0,1)$ and employ $\alpha^{-1}_{\Omega}$ to establish mappings between $B(0,1)$ and $\Omega$. Additionally, we select 200 points to encode the regions and the boundary conditions. The training dataset is structured as follows:
$\begin{split}\mathcal{T}=\{&[\phi_{200}^{1}(\Omega_{i}),\phi_{5000}^{2}(k_{\Omega_{i}}\circ\alpha^{-1}_{\Omega_{i}}),(\phi_{5000}^{2}(f_{\Omega_{i}}\circ\alpha^{-1}_{\Omega_{i}}),\phi_{200}^{3}(g_{\partial\Omega_{i}}\circ(\alpha^{-1}_{\Omega_{i}})|_{\partial B(0,1)}))],\\ &\phi_{5000}^{2}(u_{\Omega_{i}}\circ\alpha^{-1}_{\Omega_{i}})\}_{i=1}^{4500}.\end{split}$ (51)
As for the architecture, the first branch network encodes information from the input regions, represented as a tensor of size (4500, 200), with neurons [200, 500, 500, 500, 500, 1000]. The second branch network processes information from the input functions $k$ as a tensor of size (4500, 5000), with neurons [5000, 500, 500, 500, 500, 1000]. The third branch network handles information from the input function $f$ and $g$, which is represented as a tensor of size (4500, 5200), featuring only one linear layer without bias, with neurons [5200, 1000]. Additionally, the network includes a trunk network responsible for encoding the output functions within $B(0,1)$, with neurons [2, 500, 500, 500, 500, 1000]. The training parameters remain consistent with those of the previous experiment.
[…]
Figure 7: An example of prediction for fully-parameterized Poisson equation.